OpenStudy (anonymous):

How to find a point that forms a 3:5 ratio in between two points?

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (tkhunny):

3:5 ratio requires 8 pieces. 3+5 = 8 Find the distance and divide by 8. Then rethink the solution.

OpenStudy (anonymous):

@tkhunny Ok and I would do the same regardless of the ratio?

OpenStudy (tkhunny):

Yes. Find a unit that can be cut up into the right number of pieces. 2:5, well, 2+5 = 7 3:8, well, 3+8 = 11 etc.

OpenStudy (anonymous):

k thanks so much helped me understand this course a bit better!